OpenStudy (anonymous):

A stock of 12 DVD players contains two defective ones. In how many ways can one select 5 of these for inspection such that: a. Only one of the defective DVD players is included. b. Both defective DVD players are included. c. none of the defective DVD players is included.

OpenStudy (sirm3d):

a. choose one defective disc and one non-defective disc. for a defective disc, it can be selected in \(\displaystyle \binom{2}{1}\) ways. for a non-defective disc, it can be done in \(\displaystyle \binom{10}{4}\) ways.

OpenStudy (anonymous):

sirm3d is correct, but he might as well finish it off

OpenStudy (anonymous):

\[\frac{\binom{2}{1}\binom{10}{4}}{\binom{12}{5}}\] finishes it others are similar

OpenStudy (anonymous):

why divide by \[\left(\begin{matrix}12 \\ 5\end{matrix}\right)\]

OpenStudy (sirm3d):

i prefer to see his work. seeing him solve the problem is better than showing him all the solution.

OpenStudy (sirm3d):

there is no need to divide. it is not a probability question, just a counting problem.

OpenStudy (anonymous):

letter b is (2C2)(10C3) right?

OpenStudy (anonymous):

and letter c is (2C0)(10C5) right?

OpenStudy (sirm3d):

right you are!

OpenStudy (anonymous):

thanks

OpenStudy (sirm3d):

yw.